# Deep Learning Atmospheric Models Reliably Simulate Out-of-Sample Land Heat and Cold Wave Frequencies

### v1

@@ Abstract
Deep learning (DL)–based general circulation models (GCMs) are emerging as fast simulators, yet their ability to replicate extreme events outside their training range remains unknown. Here, we evaluate two such models—the hybrid Neural General Circulation Model (NGCM) and purely data-driven Deep Learning Earth System Model (DLESyM)—against a conventional high-resolution land–atmosphere model (HiRAM) in simulating land heatwaves and coldwaves. All models are forced with observed sea surface temperatures and sea ice over 1900–2020, focusing on the out-of-sample early-20th-century period (1900–1960). Both DL models generalize successfully to unseen climate conditions, broadly reproducing the frequency and spatial patterns of heatwave and cold wave events during 1900–1960 with skill comparable to HiRAM. An exception is over portions of North Asia and North America, where all models perform poorly during 1940–1960. Due to excessive temperature autocorrelation, DLESyM tends to overestimate heatwave and cold wave frequencies, whereas the physics–DL hybrid NGCM exhibits persistence more similar to HiRAM.
@@
  • Deep learning–based atmospheric models simulate 1900–1960 heatwave and coldwave frequencies with skill comparable to a physical model
  • Model architecture influences surface temperature autocorrelation, affecting the detected frequencies of simulated heat and cold waves
  • Locally poor performance over North Asia and North America suggests changes in land-surface or forcing factors other than sea-surface temperature and sea ice
@@
Heatwaves and coldwaves have major societal impacts, yet climate models often struggle to predict their occurrence. This study evaluates how well new deep learning (DL) atmospheric models simulate heatwaves and cold waves, especially when tested on early 20th-century conditions they have not encountered during training. We compare two DL-driven climate models with a traditional physics-based model, running them all with observed sea-surface temperatures from 1900 to 2020. We find that the DL models perform nearly as well as the traditional physical model in capturing heat and cold wave frequencies. Moreover, model architecture seems to influences frequency estimates by affecting the temporal persistence of surface temperature anomalies.
@@ 1 Introduction
While the models appear capable of reproducing internal atmospheric variability during the time period on which they are trained, it is less clear whether these findings hold for out-of-sample time periods (Ullrich et al., 2024). Most existing studies validate model performance under the climate conditions with abundant training data existing (e.g., Van Loon et al., 2025; Watt-Meyer et al., 2024; Kochkov et al., 2024a; Cresswell-Clay et al., 2024; Chien et al., 2025), which tends to overlap or be similar to the training set. For example, Kochkov et al. (2024a) evaluated global mean temperature from 1980 to 2020 while the model is trained on ERA5 (Hersbach et al., 2020) from 1979 to 2017. A large overlap between the training and verification periods limits assessments of model generalization. Moreover, it is uncertain to what extent these models truly capture the underlying physics versus implicitly learning the particular trends and variability of the training era. To fill this critical gap, we present an evaluation of DL-based GCMs and compare their performance to a conventional physical GCM verified against reanalysis data from 1900 to 2014. All models are forced with the same observed sea surface temperature (SST) and sea ice concentration (SIC) data, enabling a direct comparison. Here, we focus on extreme land-based temperature events, specifically heatwaves and coldwaves.
@@
Another important question concerns the influence of DL model architecture on their ability to simulate extreme events. Unlike traditional physical models, which are grounded in explicit physical theories and parameterizations that can be formally expressed mathematically, DL-based models can adopt a wide range of architectures shaped by engineering decisions, such as training efficiency and inference cost. As a result, a key question arises: how does model architecture affect the ability of models to simulate climate variability and extremes? In this study, we compare three models with differing levels of physical and data-driven components: Deep-Learning Earth System Model (DLESyM), a fully deep learning–based GCM (Cresswell-Clay et al., 2024); Neural-GCM (NGCM) a physics–DL hybrid model that blends numerical dynamics with learned subgrid processes (Kochkov et al., 2024a); and High Resolution Atmospheric Model (HiRAM), a conventional high-resolution physical model (Zhao et al., 2009). These models span a continuum from purely data-driven (DLESyM) to fully physics-based (HiRAM), allowing us to assess how the degree of explict physical constraints in the model architecture influences performance in simulating climate extremes.
@@
    How does model architecture influence the simulation of heatwaves and coldwaves?
@@ 2.1 Models
In this subsection, we describe the models used in this study. Specifically, we utilize three types of models: a purely data-driven DL-based GCM, DLESyM (Cresswell-Clay et al., 2024); a physics–DL hybrid model, NGCM (Kochkov et al., 2024a); and a traditional physics-based atmospheric GCM, HiRAM (Zhao et al., 2009). Although another purely data-driven DL-based GCM, ACE2-ERA5 (Watt-Meyer et al., 2024), was tested in this research, we do not include it in our full evaluation because its training period (1940–1995 and 2011–2019) substantially overlaps with our designated out-of-sample validation period (1900–1960), which could compromise the assessment of model generalization. Additionally, to analyze the interannual temperature variability of the AMIP experiments, we introduce a simple linear model.
@@ 2.1.4 Linear Statistical Model
To further understand the interannual temperature variability of the AMIP experiments, we train a simple linear model over the period 1980–2020, corresponding to the main training period of the deep learning models. We first apply Empirical Orthogonal Function (EOF) truncation using SACPY (e.g., Meng et al., 2023) on the HadISST dataset (Rayner et al., 2003) from 80°N to 80°S to derive the first three spatial patterns and their associated principal components (PCs) time series. We then perform a multivariate linear regression to relate the annual mean regional temperature $T(t)$to these PCs:
@@ 2.2 Datasets
20th Century Reanalysis Version 3 (20CRv3) (Slivinski et al., 2021) is a reanalysis dataset from 1836 to 2015 that we use to verify statistics of temperature extremes. It derives from assimilating surface pressure observations over ocean regions using an ensemble Kalman filter subject to prescribed sea surface temperature (SST) and sea ice concentration (SIC) boundary conditions. Here we use all 80 ensemble members to calculate the heatwave and coldwave frequencies. We note that ERA20C (Poli et al., 2016) is similar to 20CRv3, but since it only provides the ensemble mean, is insufficient for capturing event frequency distributions using the method discussed below. Additionally, ERA5 (Hersbach et al., 2020), covering 1940 to the present and incorporating a wide range of satellite and in situ observations, is used to verify our results and to compare with 20CRv3 ensemble outputs for extreme event analysis from 1960 to 2010. To further assess the reliability of 20CRv3 in capturing long-term temperature variability, we use the Berkeley Earth (BE) dataset (Rohde and Hausfather, 2020), which provides land and ocean surface temperatures from 1850 to the present. However, BE is not used directly for calculating heatwave and coldwave frequencies due to missing data between 1900 and 1960. As shown in Figure S1, the long-term temperature trends over land in regions such as Central Asia and North America differ between 20CR and BE. To further validate our experimental results, we construct a new dataset—referred to as 20CR-BE,by combining the annual mean temperatures from BE with the daily anomalies from 20CRv3 (i.e., deviations from the 20CRv3 annual mean).
@@ 3 Results
As shown in Figures 1 and 2, during the out-of-training period (1900––1960), the model-simulated spatial patterns of heatwave and coldwave frequency are broadly similar to those in 20CRv3, but with a high bias, particularly in tropical South America, central Africa, and Greenland. HiRAM exhibits smaller biases than NGCM and DLESyM, as expected from a physically based model, yet still overestimates frequency relative to 20CRv3. In the more recent period (1960–2010), HiRAM’s results align more closely with ERA5 (Figures S5 and S6). There are notable differences between 20CRv3 and ERA5, particularly over Middle Africa, Greenland and Northern South Africa. Model overestimation of heatwave and coldwave frequency is primarily associated with higher autocorrelation of daily mean surface temperature (Figure 3) as there is no large difference (mostly less than 2%) in the 90th percentile and 10th percentile threshold (Figure S3 and S4). Positive autocorrelation differences relative to 20CRv3 are spatially consistent with the frequency over-estimation in North Asia, Greenland, and central Africa. Higher autocorrelation implies greater persistence of temperature anomalies, which increases the likelihood of exceeding heatwave or coldwave consecutive days thresholds. Autocorrelation differences can be partly attributed to the AMIP SST protocol, which uses interpolated monthly SST rather than actual daily SST. As shown in Figure S11, this temporal interpolation increases autocorrelation over tropical oceans, increasing autocorrelation over nearby tropical land regions, but not globally. Despite the overestimation, the global-mean heatwave and coldwave frequencies during 1900–1960 are well correlated with 20CRv3, with correlation coefficients across all models of 0.73–0.74 for heatwaves and 0.81–0.84 for coldwaves. Replacing the annual variability of 20CRv3 with that of BE results in a small increase in correlation for all models (Figures 1k and 2k).
@@
All models show positive correlations with 20CRv3 across most regions, with particularly high correlations in the tropics as SST signal is strong in the tropical land (Figures 1e–g and 2e–g). Exceptions include North Asia and parts of the Americas, where the correlations are relatively weaker. When comparing NGCM and DLESyM with HiRAM (Figures 1h–i and 2h–i), correlations are consistently positive globally, indicating strong agreement between the three models. During the in-sample period (1961–2010), the spatial correlation with 20CRv3 improves, particularly over North Asia and the Americas, in both the physical and DL-based models (Figures S5 and S6). We attribute this improvement to better representation of interannual variability in the forcing SST during the later period, while the AMIP experiments show limited skill in reproducing such variability prior to 1960 (Figure 4). This enhancement in skill may also be partially due to a higher signal-to-noise ratio in recent decades, which facilitates the emergence of predictable patterns (Jia et al., 2016).
@@
From 1900 to 1960, the annual temperature correlations between the three models and the three observational datasets (including reanalyses) are generally weaker compared to the 1960–2010 period. Rolling correlation results (Figures 4e and 4h) further highlight periods of reduced correlation from the 1940s to the 1960s across all models. Using the simple linear model (SM) defined in section 2, we find that the decadal variability in correlation skill can be reasonably captured by the leading three modes of SST variability (recall SM is trained during 1961–2010). This suggests that the frequency variability may be related to nonstationary teleconnections to SST forcing, rather than radiative forcing from aerosols, which are absent from all models but HiRAM. Additionally, land-surface changes, such as aridity over North American during the dust bowl period of the 1930s, are not captured in AMIP experiments (Donat et al., 2016). Another potentially contributing factor may be the lower quality of SST observations in the early 20th century (e.g., Sippel et al., 2024; Chan and Huybers, 2019). For coldwaves, the low correlation over the Tibetan Plateau may be related to limited training data in this region, the relatively low reliability of reanalysis products over complex terrain (Bao et al., 2020), and for NGCM to the interpolation of temperature to 1000 hPa.
@@
By comparing results across the three models, we can draw preliminary conclusions about how model architecture may influence extreme temperature event simulation. The fully DL–based DLESyM shows the highest temperature autocorrelation, followed by the hybrid physics–DL NGCM, and then the fully physics-based HiRAM. This difference in autocorrelation results in overestimation of heatwave and coldwave frequencies in DLESyM relative to the other two models. For comparison, another fully DL-based model, ACE2-ERA5, also shows higher autocorrelation than NGCM and HiRAM, but lower than DLESyM (Figure S7).

### v2

@@ Abstract
Deep learning (DL)–based general circulation models (GCMs) are emerging as fast simulators, yet their ability to replicate extreme events outside their training range remains unknown. Here, we evaluate two such models—the hybrid Neural General Circulation Model (NGCM) and purely data-driven Deep Learning Earth System Model (DLESyM)—against a conventional high-resolution land–atmosphere model (HiRAM) in simulating land heatwaves and coldwaves. All models are forced with observed sea surface temperatures and sea ice over 1900–2020, focusing on the out-of-sample period (1900–1960). Both DL models generalize successfully to unseen climate conditions, broadly reproducing the frequency and spatial patterns of heatwave and coldwave events during 1900–1960 with skill comparable to HiRAM. An exception is over portions of North Asia and North America, where all models perform poorly during 1940–1960. Due to excessive temperature autocorrelation, DLESyM tends to overestimate heatwave and coldwave frequencies, whereas the physics–DL hybrid NGCM exhibits persistence more similar to HiRAM.
@@
  • DL-based GCMs successfully simulate out-of-sample (1900–1960) heat and cold wave frequencies with skill comparable to a physical model
  • The degree of autocorrelation in the surface temperature field for the models roughly scales with the data-driven contribution to the model
  • Locally poor performance over North Asia and North America suggests changes in the land-surface or weaker SST forcing
@@
Heatwaves and cold waves have major societal impacts, yet climate models often struggle to predict their occurrence. This study evaluates how well new deep learning (DL) atmospheric models simulate heatwaves and cold waves, especially when tested on early 20th-century conditions they have not encountered during training. We compare two DL-driven climate models with a traditional physics-based model, running them all with observed sea-surface temperatures from 1900 to 2020. We find that the DL models perform nearly as well as the traditional physical model in capturing heat and cold wave frequencies, and the probability of extreme events. Moreover, differences in model architecture influence frequency estimates by affecting the temporal persistence of surface temperature anomalies.
@@ 1 Introduction
While the models appear capable of reproducing internal atmospheric variability during the time period on which they are trained, it is less clear whether these findings hold for out-of-sample time periods (Ullrich et al., 2024). Most existing studies validate model performance under the climate conditions with abundant training data existing (e.g., Van Loon et al., 2025; Watt-Meyer et al., 2024; Kochkov et al., 2024a; Cresswell-Clay et al., 2024; Chien et al., 2025), which tends to overlap or be similar to the training set. For example, Kochkov et al. (2024a) evaluated global mean temperature from 1980 to 2020 while the model is trained on ERA5 (Hersbach et al., 2020) from 1979 to 2017. A large overlap between the training and verification periods limits assessments of model generalization. Moreover, it is uncertain to what extent these models truly capture the underlying physics versus implicitly learning the particular trends and variability of the training era. To fill this critical gap, we present an evaluation of DL-based GCMs and compare their performance to a conventional physical GCM verified against reanalysis data from 1900 to 2014. All models are forced with the same observed sea surface temperature (SST) and sea ice concentration (SIC) data, following the AMIP (Atmospheric Model Intercomparison Project) protocol (e.g., Gates et al., 1999), enabling a direct comparison. Here, we focus on extreme land-based temperature events, specifically heatwaves and coldwaves.
@@
Another important question concerns the influence of DL model architecture on their ability to simulate extreme events. Unlike traditional physical models, which are grounded in explicit physical theories and parameterizations that can be formally expressed mathematically, DL-based models can adopt a wide range of architectures shaped by engineering decisions, such as training efficiency, inference cost, and forecast accuracy. DL models used for weather and climate prediction span several major architecture families, including multilayer perceptrons (MLPs), convolutional neural networks (CNNs, e.g., U-Nets), recurrent neural networks (RNN), transformers, and physics–DL hybrids (e.g., Goodfellow et al., 2016). These model families differ in their built-in assumptions about data (inductive biases), their learning behavior during training (training dynamics), and the balance they strike between computational efficiency and physical interpretability (Goodfellow et al., 2016; Reichstein et al., 2019; Wu and Xue, 2024). As a result, a key question arises: how does model architecture affect the ability of models to simulate climate variability and extremes? In this study, we compare three models with differing levels of physical and data-driven components: Deep-Learning Earth System Model (DLESyM), a fully deep learning–based GCM (Cresswell-Clay et al., 2024); Neural-GCM (NGCM) a physics–DL hybrid model that blends numerical dynamics with learned subgrid processes (Kochkov et al., 2024a); and High Resolution Atmospheric Model (HiRAM), a conventional high-resolution physical model (Zhao et al., 2009). These models span a continuum from purely data-driven (DLESyM) to fully physics-based (HiRAM), allowing us to assess how the degree of explict physical constraints in the model architecture influences performance in simulating climate extremes.
@@
    How do purely physical, hybrid, and purely data-driven models differ in simulating heatwaves and coldwaves?
@@ 2.1 Models
In this subsection, we describe the models used in this study. Specifically, we utilize three types of models: a purely data-driven DL-based GCM, DLESyM (Cresswell-Clay et al., 2024); a physics–DL hybrid model, NGCM (Kochkov et al., 2024a); and a traditional physics-based atmospheric GCM, HiRAM (Zhao et al., 2009). Although another purely data-driven DL-based GCM, ACE2-ERA5 (Watt-Meyer et al., 2024), was tested in this research, we do not include it in our full evaluation because its training period (1940–1995 and 2011–2019) substantially overlaps with our designated out-of-sample validation period (1900–1960), which could compromise the assessment of model generalization. A brief summary of the GCMs used in our paper appears in Supplementary Table 1. Additionally, to analyze the interannual temperature variability of the AMIP experiments, we introduce a simple linear model.
@@ 2.1.4 Linear Statistical Model
To further understand the interannual temperature variability of the AMIP experiments, we train a simple linear model over the period 1980–2020, corresponding to the main training period of the deep learning models. We first apply Empirical Orthogonal Function (EOF) truncation using SACPY (e.g., Meng et al., 2023) on the HadISST dataset (Rayner et al., 2003) from 80°N to 80°S to derive the first three spatial patterns and their associated principal components (PCs) time series. We then perform a multivariate linear regression to relate the annual mean regional temperature T(t)T(t)to these PCs:
@@ 2.2 Datasets
20th Century Reanalysis Version 3 (20CRv3) (Slivinski et al., 2021) is a reanalysis dataset from 1836 to 2015 that we use to verify statistics of temperature extremes. It derives from assimilating surface pressure observations over ocean regions using an ensemble Kalman filter subject to prescribed sea surface temperature (SST) and sea ice concentration (SIC) boundary conditions. Here we use all 80 ensemble members to calculate the heatwave and coldwave frequencies. We note that ERA20C (Poli et al., 2016) is similar to 20CRv3, but since it only provides the ensemble mean, is insufficient for capturing event frequency distributions using the method discussed below. Additionally, ERA5 (Hersbach et al., 2020), covering 1940 to the present and incorporating a wide range of satellite and in situ observations, is used to verify our results and to compare with 20CRv3 ensemble outputs for extreme event analysis from 1960 to 2010. To further assess the reliability of 20CRv3 in capturing long-term temperature variability, we use the Berkeley Earth (BE) dataset (Rohde and Hausfather, 2020), which provides land and ocean surface temperatures from 1850 to the present. However, BE is not used directly for calculating heatwave and coldwave frequencies due to missing data between 1900 and 1960, except for selected locations where a direct comparison is enabled by nearly complete observation time series. As shown in Figure S1, the long-term temperature trends over land in regions such as Central Asia and North America differ between 20CR and BE. To further validate our experimental results, we construct a new dataset—referred to as 20CR-BE,by combining the annual mean temperatures from BE with the daily anomalies from 20CRv3 (i.e., deviations from the 20CRv3 annual mean).
@@ 3 Results
As shown in Figures 1 and 2, during the out-of-training period (1900––1960), the model-simulated spatial patterns of heatwave and coldwave frequency are broadly similar to those in 20CRv3, but with a high bias, particularly in tropical South America, central Africa, and Greenland. HiRAM exhibits smaller biases than NGCM and DLESyM, as expected from a physically based model, yet still overestimates frequency relative to 20CRv3. In the more recent period (1960–2010), HiRAM’s results align more closely with ERA5 (Figures S5 and S6). There are notable differences between 20CRv3 and ERA5, particularly over Equatorial Africa, Greenland and Northern South Africa. Model overestimation of heatwave and coldwave frequency is primarily associated with higher autocorrelation of daily mean surface temperature (Figure 3) as there is no large difference (mostly less than 2%) in the 90th percentile and 10th percentile threshold (Figure S3 and S4). Positive autocorrelation differences relative to 20CRv3 are spatially consistent with the frequency over-estimation in North Asia, Greenland, and central Africa. Higher autocorrelation implies greater persistence of temperature anomalies, which increases the likelihood of exceeding heatwave or coldwave consecutive days thresholds. Autocorrelation differences can be partly attributed to the AMIP SST protocol, which uses interpolated monthly SST rather than actual daily SST. As shown in Figure S11, this temporal interpolation increases autocorrelation over tropical oceans, increasing autocorrelation over nearby tropical land regions, but not globally. Despite the overestimation, the global-mean heatwave and coldwave frequencies during 1900–1960 are well correlated with 20CRv3, with correlation coefficients across all models of 0.73–0.74 for heatwaves and 0.81–0.84 for coldwaves. Replacing the annual variability of 20CRv3 with that of BE results in a small increase in correlation for all models (Figures 1k and 2k).
@@
All models show positive correlations with 20CRv3 across most regions, with particularly high correlations in the tropics as SST signal is strong in the tropical land (Figures 1e–g and 2e–g). Exceptions include North Asia and parts of the Americas, where the correlations are relatively weaker. When comparing NGCM and DLESyM with HiRAM (Figures 1h–i and 2h–i), correlations are consistently positive globally, indicating strong agreement between the three models. During the in-sample period (1961–2010), the spatial correlation with 20CRv3 improves, particularly over North Asia and the Americas, in both the physical and DL-based models (Figures S5 and S6). We also examined one simulation for the ACE2-ERA5 model and found that its performance is broadly similar to that of individual members of DLESyM, both during the in-sample period (1980–2020) and the out-of-sample period (1900–1960) (figure not shown). We attribute this improvement to better representation of interannual variability in the forcing SST during the later period, while the AMIP experiments show limited skill in reproducing such variability prior to 1960 (Figure 4). This enhancement in skill may also be partially due to a higher signal-to-noise ratio in recent decades, which facilitates the emergence of predictable patterns (Jia et al., 2016).
@@
From 1900 to 1960, the annual temperature correlations between the three models and the three observational datasets (including reanalyses) are generally weaker compared to the 1960–2010 period. Rolling correlation results (Figures 4e and 4h) further highlight periods of reduced correlation from the 1940s to the 1960s across all models. Using the simple linear model (SM) defined in section 2, we find that the decadal variability in correlation skill can be reasonably captured by the leading three modes of SST variability (recall SM is trained during 1961–2010). This suggests that the frequency variability may be related to nonstationary teleconnections to SST forcing, rather than radiative forcing from aerosols, which are absent from all models but HiRAM. Additionally, land-surface changes, such as aridity over North American during the dust bowl period of the 1930s, are not captured in AMIP experiments (Donat et al., 2016). Another potentially contributing factor may be the lower quality of SST observations in the early 20th century (e.g., Sippel et al., 2024; Chan and Huybers, 2019). Another contributing factor may be the relatively low signal-to-noise ratio in SST forcing during the mid-20th century, when global SST variability was weaker, reducing the strength of the boundary-driven signal. For coldwaves, the low correlation over the Tibetan Plateau may be related to limited training data in this region, the relatively low reliability of reanalysis products over complex terrain (Bao et al., 2020), and for NGCM to the interpolation of temperature to 1000 hPa.
@@
By comparing results across the three models, we can draw preliminary conclusions about how the degree to which physical constraints may influence extreme temperature event simulation. The fully DL–based DLESyM shows the highest temperature autocorrelation, followed by the hybrid physics–DL NGCM, and then the fully physics-based HiRAM. This difference in autocorrelation results in overestimation of heatwave and coldwave frequencies in DLESyM relative to the other two models. For comparison, another fully DL-based model, ACE2-ERA5, also shows higher autocorrelation than NGCM and HiRAM, but lower than DLESyM (Figure S7).
To assess whether out-of-sample skill primarily reflects recurrence of boundary forcing, we perform an analog-year analysis. For each out-of-sample year (1900–1960), we compute area-weighted spatial correlations between the annual-mean SST anomaly field and those from every training year (1979–2017); the in-sample year with the maximum correlation is designated the analog. We then compare, for each model, the simulated heatwave and coldwave frequency patterns between the out-of-sample year and its analog, and summarize the association between SST-pattern similarity and extremes-pattern similarity across all out-of-sample years (Figure S12). For heatwaves, the SST–extremes linkage is weak: $r=0.19$ for DLESyM, $r=0.30$ for NGCM, and $r=0.40$ for HiRAM. Coldwaves show somewhat stronger relationships, $r=0.63$ for DLESyM, $r=0.66$ for NGCM, and $r=0.50$ for HiRAM, yet the DL models (DLESyM, NGCM) exhibit correlations of similar magnitude to the physical model (HiRAM), indicating that recurrence of SST patterns alone is insufficient to explain out-of-sample performance.
Our verification analysis is based on the heatwave and coldwave definitions from Hirsch et al. (2021) and Perkins et al. (2012), which depend on the 90th and 10th percentile temperature thresholds, respectively. To complement these reanalyses-based percentile verification statistics, we now compare model simulations directly with heatwave statistics for surface temperature observations from Berkeley Earth (BE) (Rohde and Hausfather, 2020) at six capital cities—Helsinki, Paris, Washington DC, Cairo, Mexico City, and Sydney—chosen because BE daily data covers 99.9% of the out-of-sample period (1900–1960). Figure 5 shows DJF and JJA anomaly distributions (relative to 1980–2010). Despite not being trained on early-20th-century time period, the DL models reproduce the cold tails well. In particular, NGCM closely matches the DJF cold extremes in Helsinki, Paris, Washington DC, and Sydney, from 99th percentile to 99.9th percentile, even outperforming HiRAM and 20CR. The heatwave tail for these cities also shows relatively good agreement with the BE results. In contrast, for the tropical cities (Cairo and Mexico City), the DL models exhibit broader distributions and excessive probability density in the distribution tails, which may relate to higher persistence/autocorrelation in the tropics, suggesting a source of regional discrepancy that warrants further analysis.